# Non-Euclidean Erdős–Anning Theorems

### v2

@@ 3.2 Voronoi cells
As in the proof of Lemma 6, the only other possibility is that there is some other Voronoi cell $V_{j}$ also containing $p$, such that every neighborhood of $p$ contains points of $V_{j}\setminus V_{j}$. Then the shortest curves from $p$ to $s_{i}$ and to $s_{j}$ cannot follow the same ray from $p$, else $s_{i}$ would be degenerate (contradicting the assumption in the statement of the lemma) or $s_{j}$ would be degenerate (contradicting the assumption that $V_{j}\setminus V_{i}$ is nonempty). For any point $q$ past $p$ on $L$, there would exist curves of equal length from $q$ to $s_{i}$ along $L$, and from $q$ to $s_{j}$ along $L$ to $p$ and then from $p$ to $s_{j}$; but because the curve to $s_{j}$ turns at $p$ it cannot be a shortest curve as it can be shortcut near $p$, so $q$ is closer to $s_{j}$ than $s_{i}$. But this would contradict the definition of $L$ as a Voronoi segment. ∎
@@ 3.4 Erdős–Anning theorems for Riemannian manifolds
Let $s_{1}$ be any of the points, and consider the system of rays containing shortest curves to all the other points. Suppose for a contradiction that there are two rays $R_{2}$ and $R_{3}$ in this system of rays, containing shortest curves from $s_{1}$ to points $s_{2}$ and $s_{3}$ in $S$, and that these two rays are distinct and non-opposite. Then if $s_{1}$ were on a shortest path from $s_{2}$ to $s_{3}$, then there would be a shortest curve from $s_{2}$ to $s_{3}$ that follows $R_{2}$ from $s_{2}$ to $s_{1}$ and then follows $R_{2}$ from $s_{1}$ to $s_{3}$. This cannot happen, because a shortest curve that turns from one ray to another at $s_{1}$ could be shortcut by a different curve in a neighborhood of $s_{1}$. Alternatively, if $s_{2}$ were on a shortest curve from $s_{1}$ to $s_{3}$, then we could get a shortest curve from $s_{1}$ to $s_{3}$ that follows $R_{2}$ from $s_{1}$ to $s_{2}$, turns at $s_{2}$, and then continues along the $R_{3}$ from $s_{2}$ to $s_{3}$, again an impossibility. The case that $s_{3}$ is on a shortest curve from $s_{1}$ to $s_{2}$ is symmetric. So none of the three points can be the middle point of a shortest curve containing all three points. This contradiction to the assumption of the lemma implies that rays $R_{2}$ and $R_{3}$ cannot exist, and that the system of rays contains either a single ray or two opposite rays. In either case, these two rays, and therefore all points of $S$, lie on a single geodesic.
@@
If a set $S$ of points of a 2-dimensional complete Riemannian manifold of bounded genus has the property that all distances between points of $S$ are integers, then either $S$ is finite or all triples of points in $S$ are subsets of shortest curves along a single geodesic of the manifold.
@@ 4 Convex surfaces
In this section we extend our results on Riemann manifolds to non-smooth surfaces of non-negative curvature. The restriction to non-negative curvature is motivated by Example 6, which shows that even a single point of negative curvature can be an obstacle to Erdős–Anning theorems. For concreteness, in this section we define a convex surface to mean either the boundary of a convex subset of $\mathbb{R}^{3}$ with nonempty interior, or a degenerate case, the double covering of a convex subset of $\mathbb{R}^{2}$ (obtained by gluing two copies of the surface at their shared boundary). For instance, the Euclidean plane is a convex surface in this sense: it is the boundary of a halfspace. By results of Aleksei Pogorelov, a more abstract class of non-smooth convex manifolds, topologically equivalent to a sphere and with total Gaussian curvature $4\pi$, can be characterized as the bounded convex surfaces in exactly this sense, including the degenerate case [8]. We measure the distance on a convex surface as the infimum of lengths of curves on the surface. By a compactness argument, each pair of points on the surface can be connected by a curve of length equal to this distance, so this is a geodesic metric space [25], and we can apply arguments involving the Lipschitz property of distances. Every convex surface, in the sense described here, is topologically a disk (like the Euclidean plane), a sphere (like a geometric sphere, polyhedron, or double-covered topological disk), or an annulus (like the boundary of an infinite cylinder); in particular, it has Euler genus zero.
@@
Every point on a convex surface is either locally Euclidean or locally equivalent to a cone point.
@@ 4.2 Erdős–Anning theorems for convex surfaces
Same as for Theorem 1 and Theorem 4. ∎
@@
Same as for Lemma 19. ∎
@@
Same as for Theorem 5. ∎
@@
Consider the plane containing the three points; if $U$ does not intersect this plane then the points can be connected directly. Within this plane draw a line through each point, disjoint from $U$. If the three lines form a triangle or unbounded three-sided region containing $U$, with the three points on its sides, then one of its internal angles is at least $60^{\circ}$, and the curve connecting two points through this angle has length at most twice the distance between the points (Fig. 13; worst case: $U$ is an equilateral triangle and the three points are its edge midpoints). If one of the three points is not on a side of the region containing $U$, it can see the entire line through another of the points, and be connected directly to that point. ∎
@@
If $n$ points on a convex surface have diameter $D$ then some two of the points have distance $O(D/\sqrt{n})$.
@@
Let the set of points be $S$. It lies within a ball of radius $D$ in $\mathbb{R}^{3}$. Subdivide the ball by a grid of cubes, of side length $O(D/\sqrt{n})$, positioned so that none of the given points lies on a cube boundary, and discretize the convex set intersected with the ball by replacing it with the set of cubes that it intersects. A cube $C$ that contains a point of $S$ must be part of this discretized set of cubes, but it also must touch (at least at a corner) a cube that is outside the set, because if all eight cubes corner-adjacent to $C$ intersected or were within the surface, then the convex hull of the intersection points in these cubes would entirely enclose $C$, preventing it from including any part of the convex surface. Therefore, $C$ is within $O(1)$ grid positions of a boundary square of the discretized set of cubes. Any axis-parallel line can cross only two boundary squares (one in each direction), so there are $O(n)$ such squares and $O(n)$ cubes that contain points of $S$. By choosing the grid size appropriately, we can ensure that fewer than $n/2$ cubes contain points of $S$, and therefore that some cube, of side length $O(D/\sqrt{n})$, contains at least three points of $S$. We consider two cases:
@@
  • Otherwise, the convex surface is a double-covered two-dimensional convex set, two of the three points are on the same side of the double cover, and these two points can be connected directly by a segment of length $O(D/\sqrt{n})$.
@@
The example of an integer grid on a flat plane shows that the bound of Lemma 32 is tight.
@@
  • Suppose that all triples of points in $S$ that include both $p$ and $q$ belong to shortest curves. Each such point must belong either to a ray from $p$ through $q$ or a ray from $q$ through $p$, with the ray containing a shortest curve from $p$ to $q$. Further, this shortest curve (and hence these two rays) must be unique, for if there were two shortest curves between $p$ and $q$, one of them could be combined with part of a shortest curve from $p$ or $q$ to a third point $r$ to form a shortest curve that passes through $q$ or $p$ at a non-straight angle, an impossibility. Thus, there are two rays containing all remaining points in $S$, and these points must lie at integer spacing along these two rays, so there are $O(D)$ points in $S$ in total.
@@ 5 Discussion
We have shown that integer-distance point sets are finite or collinear in three broad classes of metric spaces each generalizing Euclidean distance in the plane. The next natural direction to look for a further extension of these results is in higher dimensions. Erdős and Anning state without proof that their results extend to $d$-dimensional Euclidean space [1]. However, the obvious attempt to generalize our proof to higher dimensions, by using intersection patterns of $(d+1)$-tuples of Voronoi cells, does not work for convex distance functions: there exist three-dimensional convex distance functions for which some 4-tuple of points is equidistant from arbitrarily many other points [16]. For Riemannian manifolds, the theorem does not generalize, as Example 7 shows. Nevertheless, it would be interesting to study similar questions for higher-dimensional uniform metrics, such as hyperbolic space. It would also be of interest to strengthen the bound on diameter in Theorem 6 and Theorem 9, if possible.

### v3

@@ 4.2 Voronoi cells
As in the proof of Lemma 6, the only other possibility is that there is some other Voronoi cell $V_{j}$ also containing $p$, such that every neighborhood of $p$ contains points of $V_{j}\setminus V_{j}$. Then the shortest curves from $p$ to $s_{i}$ and to $s_{j}$ cannot follow the same ray from $p$; otherwise, $s_{i}$ would be degenerate (contradicting the assumption in the statement of the lemma) or $s_{j}$ would be degenerate (contradicting the assumption that $V_{j}\setminus V_{i}$ is nonempty). For any point $q$ past $p$ on $L$, there would exist curves of equal length from $q$ to $s_{i}$ along $L$, and from $q$ to $s_{j}$ along $L$$p$ and then from $p$ to $s_{j}$; but because the curve to $s_{j}$ turns at $p$ it cannot be a shortest curve as it can be shortcut near $p$, so $q$ is closer to $s_{j}$ than $s_{i}$. But this would contradict the definition of $L$ as a Voronoi segment. ∎
@@ 4.4 Erdős–Anning theorems for Riemannian manifolds
Let $s_{1}$ be any of the points, and consider the system of rays containing shortest curves to all the other points. Suppose for a contradiction that two rays $R_{2}$ and $R_{3}$ in this system of rays contain shortest curves from $s_{1}$ to points $s_{2}$ and $s_{3}$ in $S$, and that these two rays are distinct and non-opposite. Then if $s_{1}$ were on a shortest path from $s_{2}$ to $s_{3}$, then a shortest curve from $s_{2}$ to $s_{3}$ would follow $R_{2}$ from $s_{2}$ to $s_{1}$ and then follow $R_{3}$ from $s_{1}$ to $s_{3}$. This cannot happen, because a shortest curve that turns from one ray to another at $s_{1}$ could be shortcut by a different curve in a neighborhood of $s_{1}$. Alternatively, if $s_{2}$ were on a shortest curve from $s_{1}$ to $s_{3}$, then we could get a shortest curve from $s_{1}$ to $s_{3}$ that follows $R_{2}$ from $s_{1}$ to $s_{2}$, turns at $s_{2}$, and then continues along the $R_{3}$ from $s_{2}$ to $s_{3}$, again an impossibility. The case that $s_{3}$ is on a shortest curve from $s_{1}$ to $s_{2}$ is symmetric. So none of the three points can be the middle point of a shortest curve containing all three points. This contradiction to the assumption of the lemma implies that rays $R_{2}$ and $R_{3}$ cannot exist, and that the system of rays contains either a single ray or two opposite rays. In either case, these two rays, and therefore all points of $S$, lie on a single geodesic.
@@
If a set $S$ of points of a 2-dimensional complete Riemannian manifold of bounded genus has the property that all distances between points of $S$ are integers, then either $S$ is finite or there is a single geodesic of the manifold containing all points of $S$ and connecting each pair of points by a shortest curve.
@@ 5 Convex surfaces
In this section we extend our results on Riemann manifolds to non-smooth surfaces of non-negative curvature. The restriction to non-negative curvature is motivated by Example 6, which shows that even a single cone point of negative curvature can be an obstacle to Erdős–Anning theorems. For concreteness, in this section we usually define a convex surface to mean the boundary of a convex subset of $\mathbb{R}^{3}$ with nonempty interior. However, we also allow a special case, which can be thought of as the limit of convex cylinders as the cylinder height goes to zero: the double covering of a convex subset $K$ of $\mathbb{R}^{2}$, obtained by gluing two copies of $K$ at corresponding boundary points. For instance, the Euclidean plane is a convex surface in the first of these two senses: it is the boundary of a halfspace. By results of Aleksei Pogorelov, a more abstract class of non-smooth convex manifolds, topologically equivalent to a sphere and with total Gaussian curvature $4\pi$, can be characterized as being either boundaries of convex bounded subsets of $\mathbb{R}^{3}$ or double covers of convex bounded subsets of $\mathbb{R}^{2}$ [8]. We include double covers in our definition of convex surfaces to be consistent with this result of Pogorelov, but we do not require the convex sets from which they are defined to be bounded. We measure the distance on a convex surface as the infimum of lengths of curves on the surface. By a compactness argument, each pair of points on the surface can be connected by a curve of length equal to this distance, so this is a geodesic metric space [26], and we can apply arguments involving the Lipschitz property of distances. Every convex surface, in the sense described here, is topologically a disk (like the Euclidean plane), a sphere (like a geometric sphere, polyhedron, or double-covered topological disk), or an annulus (like the boundary of an infinite cylinder).
@@
For every point $p$ on a convex surface, the disks of radius $r$ centered on that point, in the limit as $r\to 0$, approach a cone for which $p$ either has angular deficit zero ($p$ is locally Euclidean) or a well-defined positive angular deficit ($p$ is locally a cone point).
@@ 5.2 Erdős–Anning theorems for convex surfaces
Same as for Theorem 1 and Theorem 4, substituting the properties of additive Voronoi diagrams for convex surfaces in place of the corresponding properties of the diagrams for convex distance functions and Riemannian manifolds. ∎
@@
Same as for Lemma 19, again substituting the properties of Voronoi diagrams on these surfaces. ∎
@@
Same as for Theorem 5, again substituting the properties of Voronoi diagrams on these surfaces. ∎
@@
Consider the plane containing the three points; if $U$ does not intersect this plane then the points can be connected directly. Within this plane draw a line through each point, disjoint from $U$. The three lines form a triangle or unbounded three-sided region containing $U$; if all three points lie its sides, one of its internal angles is $\geq 60^{\circ}$, and the curve connecting two points through this angle has length at most twice the distance between the points (Fig. 13; worst case: $U$ is an equilateral triangle and the three points are its edge midpoints). If one of the three points is not on a side of the region containing $U$, it can see the entire line through another of the points, and be connected directly to that point. ∎
@@
If $n$ points on a convex surface have diameter $D$ then some two of the points have geodesic distance $O(D/\sqrt{n})$.
@@
[…]
@@
  • Otherwise, the convex surface is a double-covered two-dimensional convex set, two of the three points are on the same side of the double cover, and these two points can be connected directly within $C$ by a geodesic segment of length $O(D/\sqrt{n})$.
@@
The example of a square grid of $\sqrt{n}\times\sqrt{n}$ points on a flat plane, scaled to have diameter $D$, shows that Lemma 32 is tight.
@@
  • Suppose that all triples of points in $S$ that include both $p$ and $q$ belong to shortest curves. Each point in $S$ must belong either to a ray from $p$ through $q$ or a ray from $q$ through $p$, with the ray containing a shortest curve from $p$ to $q$. Further, this shortest curve (and hence these two rays) must be unique, for if there were two shortest curves between $p$ and $q$, one of them could be combined with part of a shortest curve from $p$ or $q$ to a third point $r$ to form a shortest curve that passes through $q$ or $p$ at a non-straight angle, an impossibility. Thus, two rays contain all remaining points in $S$, and these points lie at integer spacing along these two rays, giving $O(D)$ points in $S$ in total.
@@ 6 Discussion
We have shown that integer-distance point sets are finite or collinear in three broad classes of metric spaces each generalizing Euclidean distance in the plane. The next natural direction to look for a further extension of these results is in higher dimensions. Erdős and Anning state without proof that their results extend to $d$-dimensional Euclidean space [1]. However, the obvious attempt to generalize our proof to higher dimensions, by using intersection patterns of $(d+1)$-tuples of Voronoi cells, does not work for convex distance functions: there exist three-dimensional convex distance functions for which some 4-tuple of points is equidistant from arbitrarily many other points [17]. For Riemannian manifolds, the theorem does not generalize, as Example 7 shows. Nevertheless, it would be interesting to study similar questions for higher-dimensional uniform metrics, such as hyperbolic space. It would also be of interest to strengthen the bound on diameter in Theorem 6 and Theorem 9.